# Principal stratification with $U$-statistics under principal ignorability

### v4

@@ 2.3 Identification assumptions
Assumption 1 rules out confounding between the treatment assignment and intermediate variable as well as final outcomes conditional on observed pre-treatment covariates. This assumption holds by design in randomized experiments. For observational studies, Assumption 1 can be satisfied if all confounders (common causes) affecting the treatment assignment and the intermediate variable, and those affecting the treatment assignment and the final outcome are measured and included in $X$.
@@
Assumption 2 states that the treatment assignment only affects the intermediate outcome in a non-negative (or non-positive if assume $D_{i}(1)\leq D_{i}(0)$) direction and restricts the number of possible strata by excluding the harmed or defiers stratum $S=01$ (or $S=10$). For example, when $D$ represents the survival status, monotonicity holds when the treatment is expected not to harm the survival of each unit. When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to active treatment.
@@
Given the same values of pre-treatment covariates for a pair of units ($X_{1}=x_{1}$ and $X_{2}=x_{2}$), the expected contrast function is exchangeable across certain pairs of principal strata. That is,
$\begin{split}&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2},X_{1}=x_{1},X_{2}=x_{2}]\\ =~&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1}^{\prime},S_{2}=s_{2}^{\prime},X_{1}=x_{1},X_{2}=x_{2}]\end{split}$ (6)
@@
Intuitively, Assumption 3 rules out unmeasured confounding between principal stratum membership and the final outcome, thereby ensuring that the PGCE estimands can be identified from the observed data. Under Assumptions 1 - 3, for instance, the conditional expectations within group $\mathcal{S}_{10}$ all reduce to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$. In other words, within the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the various combinations of possible latent strata including $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$ are exchangeable with respect to the expected contrast. Similarly, the conditional expected contrasts within $\mathcal{S}_{11}$, $\mathcal{S}_{00}$, and $\mathcal{S}_{01}$ are all equal to the estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$, respectively (see Table 1(a) for a concrete mapping between the observed data strata and the latent principal strata). That is, with Assumption 3, each one of the observed subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only (Table 1(b)).
@@
Assumption 3 can be viewed as a generalization of the unitwise mean principal ignorability assumption in Jiang et al. (2022), since it reduces to that form in the linear contrast case h(u,v)=uvh(u,v)=u-v. For nonlinear contrasts, however, expectations cannot be interchanged within the contrast function, and thus there is no clear nesting relationship between the pairwise and unitwise versions of mean principal ignorability. Consider the win estimands under noncompliance (Example 2) with the contrast function $h(u,v)=\mathbbm{1}(u>v)$. Applying (LABEL:eq:pi-equation) to the first two pairs in $\mathcal{S}_{10}$ yields $P\{Y_{1}(1)>Y_{2}(0)|S_{1}=10,S_{2}=10,X_{1}=x_{1},X_{2}=x_{2}\}=P\{Y_{1}(1)>Y_{2}(0)|S_{1}=11,S_{2}=10,X_{1}=x_{1},X_{2}=x_{2}\}$. This assumption states that, conditional on the same pairwise covariate profiles, the probability that a treated individual has a more favorable outcome than a control individual is the same whether the pair consists of two compliers or of one complier and one always-taker. In other words, there do not exist additional unmeasured factors influencing the conditional probability of winning among pairs in $\mathcal{S}_{10}$. This is neither stronger nor weaker than the unitwise mean principal ignorability assumption, which requires, for example, $\mathbb{E}\{Y_{1}(1)|S_{1}=10,X_{1}=x_{1}\}=\mathbb{E}\{Y_{1}(1)|S_{1}=11,X_{1}=x_{1}\}$.
@@ 11 $0$ (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
Finally, we distinguish Assumptions 1 and 2 from Assumption 3. The first two are stated at the individual level, but since the observed data are assumed to be i.i.d., they extend naturally to pairs. For example, individual-level treatment ignorability implies that $(Z_{1},Z_{2})$ is jointly independent of the potential outcomes and intermediate variables given $(X_{1},X_{2})$. By contrast, Assumption 3 must be formulated at the pairwise level in its minimally sufficient form, because identification with nonlinear contrasts requires linking expectations across observed and latent strata pairs as in Table 1. Assumptions 2 and 3 are not testable from the observed data alone. We therefore complement our main results with sensitivity analyses under violations of these assumptions, provided in Section 4.4 and Section S7 of the Online Supplement.
@@ 3 Nonparametric identification
3.1 Identification of the PGCE estimands
To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution. We consider both the observational and experimental studies and define the treatment propensity score by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin, 1983). In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu, 2016), for $s\in\{10,00,11\}$ under Assumption 2. Furthermore, we use $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$ to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively. Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\}$, for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Below, we provide two examples of the pairwise outcome mean models for specific contrast functions.
@@ 3.1 Identification of the PGCE estimands
For studying the probabilistic index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered such that $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$. In this case, one can derive the pairwise outcome mean as
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\Phi\left\{\frac{f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2})}{\sqrt{2}\sigma}\right\}.$ (7)
@@
Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$. For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ (8)
$\quad=\left\{\sum_{q=2}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2}),\sum_{q=1}^{Q-1}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2})\right\},$
where $p_{zq}(D,X)$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $X$, and $Q(\geq 2)$ is the largest category of the ordinal outcome.
For estimating the principal scores, we first define the conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$. Then, under Assumption 2, the principal scores can be expressed as $e_{10}(X)=p_{1}(X)-p_{0}(X)$, $e_{00}(X)=1-p_{1}(X)$, and $e_{11}(X)=p_{0}(X)$. The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$. Based on these elements, we establish three alternative nonparametric identification formulas for each PGCE estimand, as shown in Theorem S2 of the Online Supplement. Each formula requires specifying only two of the three models—the treatment propensity score, the principal score, and the pairwise outcome mean—rather than all three simultaneously. Furthermore, Theorem S1 of the Online Supplement presents the associated balancing conditions for any pairwise covariate functions. These balancing conditions provide the theoretical foundation for model checking of the treatment propensity and principal score models without peeking at the outcome data.
@@ 3.2 Estimators based on nonparametric identification formulas
For point estimation, a common practice for implementing the nonparametric identification formulas is through specifying parametric regression models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$. We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$. To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment propensity, the principal score, and the outcome mean models are correctly specified, respectively. That is, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Additionally, we use the logic notation ‘+’ to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts and expectations with empirical averages. For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$. Define the following quantity for any function $f(Y,D,X)$:
$\begin{aligned} \psi_{f(Y_{z},D_{z},X)}(\mathcal{O})&=\frac{\mathbbm{1}(Z=z)[f(Y,D,X)-\mathbb{E}\{f(Y,D,X)|X,Z=z\}]}{\mathbb{P}(Z=z|X)}\\ &\qquad+\mathbb{E}\{f(Y,D,X)|X,Z=z\}\end{aligned}.$ (9)
One example of (9) is $f(Y_{z},D_{z},X)=D_{z}$, in which case we obtain the following function:
$\psi_{D_{z}}(\mathcal{O})=\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X)\}}{\mathbb{P}(Z=z|X)}+p_{z}(X).$ (10)
(10) is useful for the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome:
@@
For any general constrast function $h$, the plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart, 1998, §12). For example, the identification results in Theorem S2 of the Online Supplement motivate the following estimators for $\tau_{h}^{10}$:
$\widehat{\tau}_{h,\pi+p}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right],$
$\widehat{\tau}_{h,\pi+\mu}^{10}$ $=\mathbb{U}_{n}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1};\widehat{\alpha})-D_{1}(1-Z_{1})/\{1-\pi(X_{1};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\right$.

### v5

@@ 2.2 Identification assumptions
Assumption 1 rules out confounding between the treatment assignment and intermediate variable as well as final outcomes conditional on observed pre-treatment covariates. This assumption holds by design in randomized experiments. For observational studies, Assumption 1 can be satisfied if all confounders (common causes) affecting the treatment assignment and the intermediate variable, and those affecting the treatment assignment and the final outcome are measured and included in $\bm{X}$.
@@
Assumption 2 states that the treatment assignment only affects the intermediate outcome in a non-negative (or non-positive if assume $D_{i}(1)\leq D_{i}(0)$) direction and restricts the number of possible strata by excluding the harmed or defiers stratum $S=01$ (or $S=10$). For example, when $D$ represents the survival status, monotonicity holds if the treatment is not expected to harm the survival of each unit. When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to the active treatment.
@@
Given the same values of pre-treatment covariates for a pair of units ($\bm{X}_{1}=\bm{x}_{1}$ and $\bm{X}_{2}=\bm{x}_{2}$), the expected contrast function is exchangeable across certain pairs of principal strata. That is,
$\begin{split}&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2},\bm{X}_{1}=\bm{x}_{1},\bm{X}_{2}=\bm{x}_{2}]\\ =~&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1}^{\prime},S_{2}=s_{2}^{\prime},\bm{X}_{1}=\bm{x}_{1},\bm{X}_{2}=\bm{x}_{2}]\end{split}$ (6)
@@
Assumption 3 rules out unmeasured confounding between principal stratum membership and the final outcome, thereby ensuring that the PGCE estimands can be identified from the observed data. Under Assumptions 1 - 3, for instance, the conditional expectations within group $\mathcal{S}_{10}$ all reduce to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,\bm{X}_{1},\bm{X}_{2}]$. In other words, within the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the various combinations of possible latent strata including $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$ are exchangeable with respect to the expected contrast. Similarly, the conditional expected contrasts within $\mathcal{S}_{11}$, $\mathcal{S}_{00}$, and $\mathcal{S}_{01}$ are all equal to the estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,\bm{X}_{1},\bm{X}_{2}]$, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,\bm{X}_{1},\bm{X}_{2}]$, and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1,\bm{X}_{1},\bm{X}_{2}]$, respectively (see Table 1(a) for a mapping between the observed data strata and the latent principal strata). That is, with Assumption 3, each one of the observed subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=1)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only (Table 1(b)).
@@
Assumption 3 is a generalization of the unitwise mean principal ignorability assumption in Jiang2022, since it reduces to that form in the linear contrast case h(u,v)=uvh(u,v)=u-v. For nonlinear contrasts, however, expectations cannot be interchanged within the contrast function, and thus there is no clear nesting relationship between the pairwise and unitwise versions of mean principal ignorability. Consider the win estimands under noncompliance (Example 2) with the contrast function $h(u,v)=\mathbbm{1}(u>v)$. Applying (LABEL:eq:pi-equation) to the first two pairs in $\mathcal{S}_{10}$ yields $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=10,S_{2}=10,\bm{X}_{1}=\bm{x}_{1},\bm{X}_{2}=\bm{x}_{2}\}=\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=11,S_{2}=10,\bm{X}_{1}=\bm{x}_{1},\bm{X}_{2}=\bm{x}_{2}\}$. This assumption states that, conditional on the same pairwise covariate profiles, the probability that a treated individual has a more favorable outcome than a control individual is the same whether the pair consists of two compliers or of one complier and one always-taker. In other words, there do not exist additional unmeasured factors influencing the conditional probability of winning among pairs in $\mathcal{S}_{10}$. This is neither stronger nor weaker than the unitwise mean principal ignorability assumption, which requires, for example, $\mathbb{E}\{Y_{1}(1)|S_{1}=10,\bm{X}_{1}=\bm{x}_{1}\}=\mathbb{E}\{Y_{1}(1)|S_{1}=11,\bm{X}_{1}=\bm{x}_{1}\}$.
@@ 11 $0$ (a) $\{Z_{1}=1,D_{1}=1\}$ & $\{Z_{2}=0,D_{2}=0\}$
Finally, we distinguish Assumptions 1 and 2 from Assumption 3. The first two are stated at the individual level, but since the observed data are assumed to be i.i.d., they extend naturally to pairs. For example, individual-level treatment ignorability implies that $(Z_{1},Z_{2})$ is jointly independent of the potential outcomes and intermediate variables given $(\bm{X}_{1},\bm{X}_{2})$. By contrast, Assumption 3 must be formulated at the pair level in its minimally sufficient form, because identification with nonlinear contrasts requires linking expectations across observed and latent strata pairs as in Table 1. Assumptions 2 and 3 are not testable from the observed data alone. We therefore complement our main results with sensitivity analyses under violations of these assumptions, provided in Section 4.4 and Section S7 of the Online Supplement.
@@ 3 Nonparametric identification
3.1 Identification of the PGCEs
To obtain nonparametric identification formulas for the PGCEs, we start by defining components of the observed data distribution. We consider both observational and experimental studies and define the treatment propensity score as $\pi(\bm{X})=\mathbb{P}(Z=1|\bm{X})$, the probability that a participant receives the treatment conditional on pre-treatment covariates (Rosenbaum1983). In addition, the principal score, $e_{s}(\bm{X})=\mathbb{P}(S=s|\bm{X})$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding2016), for $s\in\{10,00,11\}$ under Assumption 2. Furthermore, we define $\pi=\mathbb{E}\{\pi(\bm{X})\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(\bm{X})\}=\mathbb{P}(S=s)$. Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as $\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},\bm{X}_{1},\bm{X}_{2}\}$, for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Below, we provide two examples of the pairwise outcome mean models for specific contrast functions.
@@ 3.1 Identification of the PGCEs
To study the probabilistic index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered such that $Y(z)|D,\bm{X}\sim\mathcal{N}(f_{z}(D,\bm{X}),\sigma^{2})$. In this case, one can derive the pairwise outcome mean as
$\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2})=\Phi\left\{\frac{f_{z_{1}}(d_{1},\bm{X}_{1})-f_{z_{2}}(d_{2},\bm{X}_{2})}{\sqrt{2}\sigma}\right\}.$ (7)
@@
Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$. For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $\bm{X}$, Mao2018 has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
$\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2})$ (8)
$\quad=\left\{\sum_{q=2}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},\bm{X}_{1})p_{z_{2}q^{\prime}}(d_{2},\bm{X}_{2}),\sum_{q=1}^{Q-1}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},\bm{X}_{1})p_{z_{2}q^{\prime}}(d_{2},\bm{X}_{2})\right\},$
where $p_{zq}(D,\bm{X})$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $\bm{X}$, and $Q(\geq 2)$ is the largest category of the ordinal outcome.
For estimating the principal scores, we first define the conditional probability, $p_{z}(\bm{X})=\mathbb{P}(D=1|Z=z,\bm{X})$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(\bm{X})\}=\mathbb{P}(D=1|Z=z)$. Then, under Assumption 2, the principal scores can be expressed as $e_{10}(\bm{X})=p_{1}(\bm{X})-p_{0}(\bm{X})$, $e_{00}(\bm{X})=1-p_{1}(\bm{X})$, and $e_{11}(\bm{X})=p_{0}(\bm{X})$. The population fractions of the three principal strata are $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(\bm{X}),p_{1}(\bm{X})\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(\bm{X}),e_{00}(\bm{X}),e_{11}(\bm{X})\}$. Based on these elements, we establish three alternative nonparametric identification formulas for each PGCE estimand, given in Theorem S2 of the Online Supplement. Each formula requires specifying two of the three models—the treatment propensity score, the principal score, and the pairwise outcome mean—rather than all three simultaneously. Furthermore, Theorem S1 of the Online Supplement presents the associated covariate balancing conditions. These balancing conditions provide the foundation for model checking of the treatment propensity and principal score models without peeking at the outcome data.
@@ 3.2 Estimators based on nonparametric identification formulas
For point estimation, a common approach is to specify parametric regression models for nuisance functions, i.e., $\pi(\bm{X};\alpha)$, $p_{z}(\bm{X};\beta)$ or $e_{s}(\bm{X};\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$. We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$. To aid further discussions of the statistical properties of estimators for the PGCEs, we introduce notation $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote the conditions under which the treatment propensity, principal score, and outcome mean models are correctly specified, respectively. That is, under $\mathcal{M}_{\pi}$, we have $\pi(\bm{X};\alpha^{*})=\pi(\bm{X})$; under $\mathcal{M}_{p}$, we have $p_{z}(\bm{X};\beta^{*})=p_{z}(\bm{X})$ and $e_{s}(\bm{X};\beta^{*})=e_{s}(\bm{X})$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(\bm{X}_{1},\bm{X}_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Additionally, we use the logic notation ‘+’ to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts and expectations with empirical averages. For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$. Define the following quantity for any function $f(Y,D,\bm{X})$:
$\begin{aligned} \psi_{f(Y_{z},D_{z},\bm{X})}(\mathcal{O})&=\frac{\mathbbm{1}(Z=z)[f(Y,D,\bm{X})-\mathbb{E}\{f(Y,D,\bm{X})|\bm{X},Z=z\}]}{\mathbb{P}(Z=z|\bm{X})}\\ &\qquad+\mathbb{E}\{f(Y,D,\bm{X})|\bm{X},Z=z\}\end{aligned}.$ (9)
One example of (9) is $f(Y_{z},D_{z},\bm{X})=D_{z}$, in which case we obtain the following function:
$\psi_{D_{z}}(\mathcal{O})=\frac{\mathbbm{1}(Z=z)\{D-p_{z}(\bm{X})\}}{\mathbb{P}(Z=z|\bm{X})}+p_{z}(\bm{X}).$ (10)
Equation (10) is useful for the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome:
@@
For any general contrast function $h$, the plug-in estimators for the PGCEs take the form of U-statistics (vandervaart1998, §12). For example, the identification results in Theorem S2 of the Online Supplement motivate the following estimators for $\tau_{h}^{10}$:
$\widehat{\tau}_{h,\pi+p}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(\bm{X}_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(\bm{X}_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(\bm{X}_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(\bm{X}_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{0}(\bm{X}_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(\bm{X}_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right],$
$\widehat{\tau}_{h,\pi+\mu}^{10}$ $=\mathbb{U}_{n}\left(\left[\frac{D_{1}Z_{1}/\pi(\bm{X}_{1};\widehat{\alpha})-D_{1}(1-Z_{1})/\{1-\pi(\bm{X}_{1};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\right$.